OpenStudy (anonymous):

Simplify. Please show all work all your work. 3c2+4c+4-(c-6c2-8)

OpenStudy (anonymous):

Is it $$ 3c^2+4c+4-(c-6c^2-8) $$ ??

jhonyy9 (jhonyy9):

3c2 +4c+4-c+6c2+8=9c2 +3c +12=3(3c2+c+4)